OpenStudy (anonymous):

suppose a population of 250 crickets doubles in size every 3 months. How many crickets will there be after 2 years ?

OpenStudy (lieutenantgeneral):

24 / 3 = 8 250 * 2 = 500 500 * 24 = 12000

OpenStudy (anonymous):

how many months in 2 years?

OpenStudy (anonymous):

the above answer is not right, btw

OpenStudy (anonymous):

24 @satellite73

OpenStudy (anonymous):

ok and \(24\div 3=8\) sor sure but that means it doubles 8 times !

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

instead of doubling 8 times, by writing \[250,500,1000,...\] it is easier to compute \[250\times 2^8\]

OpenStudy (anonymous):

that was right.